# Object as a Service: Simplifying Cloud-Native Development through Serverless Object Abstraction

### v1

@@ IV-B Overview of the Oparaca Architecture
Figure 4: A bird-eye view to the architecture of Oparaca. Dashed lines show the workflow of actions initiated by the developer to define classes and objects, and solid lines show them for the end-user to make use of objects and invoke their functions either synchronously (red arrows) or asynchronously (blue ones).
@@
  • Object Module serves as the cornerstone of Oparaca that has two main objectives: (a) providing the “object access interface” for the user application to access an object(s); and (b) offering the object abstraction while transparently handle function invocation and state manipulation.
  • FaaS Engine is the underlying execution plane of Oparaca, which can be any of the existing FaaS systems (e.g., Knative).
  • Data Management Module that handles the objects’ state.
@@ IV-C Class Module
To define a new class and its functions in Oparaca, the developer must bundle them as one package and register them to the Package Manager, shown in Figure 4. Upon successful package validation by the Package Manager, the Class Runtime Manager (termed CRM for brevity) performs the class registration process that includes two operations:
(a) Informing the Object Module about the new/updated class. Upon receiving a class registration request, the Object Module creates a handler instance to be prepared for handling object invocation requests. We elaborate on this process in Section IV-D.
(b) Registering the custom functions of the new class in the FaaS engine, so that they can be invoked at a later time. Recall that we aim to make Oparaca agnostic from the underlying FaaS engine. Therefore, the function registration process should be doable for various FaaS engines. To that end, we design Oparaca to be able to host a dedicated CRM for each FaaS engine. Therefore, a new FaaS engine can be integrated into Oparaca by simply plugging its dedicated CRM into the system. Upon receiving the function registration event, only the specific CRM associated with the configured FaaS engine processes the event, which includes translating the developer-defined function configuration into the FaaS-engine-specific format and then forwarding it to the associated FaaS engine (e.g., Knative). Consequently, the underlying FaaS Engine creates the actual function runtime to be invoked by the Object Module.
@@ IV-D Object Module and FaaS Engine
Figure 5: The cluster of Invokers replicates and distributes object data across the cluster memory via consistent hashing technique and in-memory data grid (IMDG) to reduce the data transfer overhead. The invoker offloads the invocation to a corresponding FaaS engine function that must pass through the reverse proxy.
@@
Recall that OaaS needs to support three types of functions: built-in, custom, and dataflow. Unlike built-in functions (those provided by the platform) and dataflow functions (those to handle developer-defined DAGs) that can be executed without the direct need of the FaaS engine, custom functions need to execute the developer-provided code on the FaaS engine. Thus, Oparaca requires a mechanism to utilize the FaaS engine to execute the custom function code while allowing it to access the object state transparently and with the minimum data transfer overhead. Needless to say, this mechanism also maintains the separation between the Object Module and the FaaS engine.
@@
To fulfill the aforementioned expectations, in the Object Module, we design the object invocation mechanism to be cognizant of the type of state (structured vs unstructured) and manage it such that the data transfer overhead needed for the invocation is minimized. We develop a hybrid approach that leverages the “pure function” technique to handle invocations with structured data and the “datastore abstraction” technique to handle cases with unstructured data. The rationale of this design choice is that the unstructured state (i.e., BLOB) is usually large and expensive to transfer, hence, to maintain efficiency, the FaaS engine should retrieve the state directly from the object storage (e.g., S3) in a lazy (on-demand) manner. This differs from the structured state for which we include the state as an input argument to maintain a clear separation between the Object Module and the FaaS engine and let the FaaS engine maintain its statelessness.
@@
In the Oparaca architecture (Figure 4), the described invocation and state handling mechanism is handled by the Invoker component. In particular, to offload the object invocation to the FaaS engine, Invoker bundles the request and the related structured object data as a “task”, as described in the next part, and passes it to the associated FaaS engine for execution.
IV-D1 Task Generation in the Invoker Component
@@
To further reduce the data transfer overhead of providing the object abstraction in the task generation process, we design Invokers to maintain the object data (i.e., state and metadata) in a distributed hash table [19], thereby reducing the cost of data transfer in a scalable manner. As shown in Figure 5, we equip each Invoker instance with an embedded in-memory data grid (IMDG) [43]. IMDG partitions the entire DHT into multiple segments and distributes them across Invoker instances. The invoker with IMDG determines the segment for a given object by consistent hashing of the object ID and assigns the object data to the selected segment. Similarly, to retrieve the object data, IMDG determines the object data location and then fetches it from the owner of the segment in at most one hop. Upon receiving a function call, the Invoker bundles the invocation request and associated object data into the task and offloads it to be executed on the FaaS engine.
@@ IV-D3 Task Completion
After the FaaS engine completes the task, it sends the task completion data to the Invoker to update the state. If the function reports a failed task, the state remains unchanged. Otherwise, the Invoker updates the object data in IMDG and then writes it to the persistent database immediately or asynchronously. If an invocation involves both structured and unstructured states, we use pure and datastore techniques together, which can potentially lead to ”state inconsistency challenges”. We address this challenge in section IV-F.
@@ IV-D4 Synchronous and Asynchronous Invocation
As mentioned in Section IV-A and shown in Figure 5, we designed Oparaca to offer two modes of function invocation: synchronous and asynchronous. In the synchronous mode, the function is executed immediately upon invocation and returns the result to the caller. Meanwhile, in the asynchronous mode, the invocation ID is provided to the caller as a reference, so they can later check the invocation result. The request is placed into the message broker to be reliably processed at a later time. To accommodate both modes, the Invoker utilizes the handler instance to accept the invocation request for either the REST API (synchronous) or the message broker (asynchronous). Subsequently, the handler instance forwards the request to be processed in the same way by the other part of the Invoker.
@@ IV-F2 Resilience against race condition
Race conditions in Oparaca can occur when multiple invocations modify the same object data simultaneously, resulting in potential data inconsistency. One way to prevent this issue is by using database transactions; however, this method lacks abstraction as it allows direct function code access to the database and is tightly dependent on the type of database. An alternative approach to avoiding race conditions is the cluster-wide pessimistic locking mechanism to synchronize the locking state for all invokers. Nevertheless, this approach necessitates additional network communication to coordinate the locking state, which can lead to scalability issues. Alternatively, we develop an improved version of this mechanism, called ”localized locking,” which relies on consistent hashing to direct the invocation request to the invoker that owns the primary copy of the targeted object data. Each invoker will only need to lock the object locally without additional network communication, making it more scalable than the cluster-wide version. Additionally, our localized locking approach guarantees that requests to the same object are executed in the arrival order, which is necessary in certain use cases where order matters, such as seat reservations. This is difficult to achieve with cluster-wide locking.
@@ IV-F3 Failure recovery in Oparaca
However, the message broker cannot completely address the third problem. That is, the Invoker can process the same invocation request more than once when the message broker has not acknowledged the completed one before the system failure occurs. We prevent this problem by tracking the offset number of the last processed request and adding it to each object metadata. In this manner, before processing an invocation request, Invoker checks the offset number of the target object to see if it is lower than the offset number of an incoming request. When the condition is met, the Invoker can detect it has not been processed and perform the normal operation. Otherwise, it must be skipped to avoid reprocessing.
@@ IV-G Dataflow Abstraction in Oparaca
To offer a higher level abstraction to declare a workflow of invocations, Oparaca provides the dataflow (as opposed to the standard workflow) abstraction as a built-in feature to enable developers to define the invocation steps in the form of a directed acyclic graph (DAG). In every step, the developer can declare the output of each invocation as a temporary variable within the workflow. Then, the next invocation can use the temporary variables from previous steps as the input or target to call the function. Upon the developer registering a dataflow function, Invoker constructs the DAG by having the invocation step as the edge and the objects as nodes.
When a request for dataflow execution is received, one of the Invokers takes on the role of orchestrator, similar to the orchestrator pattern [34] in microservices. It breaks down the dataflow into multiple lower-level invocations and forwards them based on the topological order of DAG. Using consistent hashing, the invoker can determine the address of the target object and send the request directly to another Invoker that holds the target object. When each step is completed, the orchestrator keeps track of the intermediate dataflow state to transparently operate the data exchange between invocation steps. With the orchestrator pattern, the dataflow control logic is centralized into a single invoker, simplifying the management, monitoring, and error-handling implementation.
When using the orchestrator pattern, the exact-once guarantee may be compromised because the object data is stored separately from the dataflow state. If the guarantee is needed, Oparaca allows developers to mitigate this problem by flagging all invocation steps on as immutable. Upon handling the dataflow request, Oparaca can generate the output ID in advance for each step, making each step of dataflow execution idempotent and safely re-executable.
@@ V-1 Security
Certain security measures are put in place in Oparaca to strengthen it against potential attacks. The first measure is to reduce the attacking surface by limiting the necessary outbound traffic from the function runtime. As the function container only requires access to the storage adapter and object storage, the traffic policy can be configured to block outbound traffic except for the Storage Adapter and object storage. The second measure is to avoid reusing secret tokens. To prevent the function container from accessing out-of-context data via analyzing the URL path, we use the presigned URL mechanism for object storage. Thus, object storage in Oparaca is more secure than in FaaS, where the same secret key is used for every request. To secure the storage adapter, we can make the Invoker generate a unique secret token for each task, and every request for the storage adapter must be authenticated via the secret token.
V-2 Cold Start
The developer functions and the Oparaca components can benefit from scale-to-zero to reduce the cost when there is no usage. However, this has the side-effect of more cold starts. Since Oparaca components are shared across functions, we can effectively keep it warm to eliminate the additional cold start impact. In such a case, the cold start performance entirely depends on the underlying serverless execution engine.
@@ VI Performance Evaluation
Figure 7: The average execution time of functions for objects with various state sizes in synchronous and asynchronous invocations. Two versions of Oparaca are examined: the full version and without URL-redirection version (oprc-relay). We also capture the time used by the internal Knative part in both Oparaca versions and show them with suffix -exec and plot it in the same bar as its Oparaca version.
@@ VI-A Experimental Setup
We deploy the Oparaca platform on 4 machines of Chameleon Cloud [26], each with 2 sockets of 24-Core Intel(R) Xeon(R) Gold 6240R CPU processors that collectively have 192 cores, 768 GB memory, and SSD SATA storage. We set up the Kubernetes cluster, which includes 15 VMs with 16 vCPUs and 32 GB memory. We made another 2 VMs for the S3-compatible storage (Minio [23]) for unstructured data and ArangoDB ([21]) for structured data. Oparaca is implemented using Java [6] with Infinispan [24] for IMDG.
@@
We used Gatling [11] for load generation and implemented three applications to serve as the workload. First is the video transcoding function using FFmpeg [42], which is a CPU-intensive application. The second is a text concatenation function that concatenates the content of a text file (state) with an input string. This function represents a highly IO-incentive workload. Third is the JSON update function, which has only structured data as a JSON and is used for randomly putting key-value pairs into the JSON state data. The rest of the workload characteristics are specific to each experiment and are explained in the respective sections. All three functions are implemented in Python language [39].
@@ VI-B Analyzing the Imposed Overhead of Oparaca
For the transcoding function (Figures 8a and 8d), at the low concurrency levels ($<$ 80 invocations), Oparaca has average response times higher than Knative, but for the higher concurrency levels, the response time of Knative grows faster than Oparaca due to computing resource limitations. Oparaca doesn’t need to fetch video file metadata, giving it an edge at high concurrency. In the concatenation function (Figures 8b and 8e), however, this phenomenon does not happen. The difference is that text concatenation is IO-intensive and desires high network bandwidth. The overhead of unstructured data access overwhelms the performance gain from eliminating structured data fetching.
@@ VI-C Scalability of the Oparaca Platform
According to Figure 9a, the speedup of Knative plateaus after reaching 6 VMs. After a thorough investigation, we realized that this plateau is attributed to the database write operation throughput bottleneck. Conversely, Oparaca exhibits the potential for higher speedup enhancement due to its reliance on the distributed in-memory hash table to consolidate data for batch write operations. This approach can boost maximum throughput by up to 3.27x when comparing oprc-bypass with knative.
@@ VI-D Performance of Localized Locking
In order to analyze the effectiveness of localized locking, we created an alternative version of Oparaca called oprc-cl that is cluster-wide locking. We then conducted an evaluation using a cluster of 12 Invokers and increasing the request arrival rate to measure the overhead of the locking mechanism. To generate requests involving the locking mechanism, we created multiple requests targeting the same object. From Figure 10, the overhead of localized locking remains mostly constant, whereas the overhead of cluster-wide locking rises with higher request rates. However, the cluster-wide version does not display this behavior, as the network communication overhead limits the throughput and hinders high invocation rates.

### v2

@@ IV-B Overview of the Oparaca Architecture
Figure 4: A bird-eye view of the Oparaca architecture. Dashed lines show actions of the developer defining classes and objects, and solid lines show actions using objects and invoking functions.
@@
  • Object Module serves as the cornerstone of Oparaca that has two main objectives: (a) providing the “object access interface” for the user application to access an object(s); and (b) offering the object abstraction while transparently handling function invocation and state manipulation.
  • FaaS Engine is the underlying execution engine of Oparaca, which can be any existing FaaS system (e.g., Knative).
  • Data Management Module is to manage object data persistence via employing database (e.g., document database) and object storage (e.g., S3-compatible storage). To bind these storages to the functions, the Invoker abstracts data access for structured data, while the Storage Adapter is employed to handle access to unstructured data in the object storage
@@ IV-C Class Module
To define a new class and its functions in Oparaca, the developer defines them as a package definition and registers it to the Package Manager, shown in Figure 4. Upon successful package validation by the Package Manager, the Class Runtime Manager (termed CRM for brevity) performs the class registration process that includes two operations:
(a) Informing the Object Module about the new/updated class. Upon receiving a class registration, the Object Module creates a handler instance to be prepared for handling object invocation. We elaborate on this process in Section IV-D.
(b) Registering the custom functions of the new class in the FaaS engine for future invocation. Recall that we aim to make Oparaca agnostic from the underlying FaaS engine. We design Oparaca to host a dedicated CRM for each FaaS engine. Accordingly, a new FaaS engine can be integrated into Oparaca by simply plugging its dedicated CRM into the system. When a function registration event occurs, the corresponding CRM processes this event by translating the function configuration into the specific format for that engine (e.g., Knative) and forwards it. Consequently, the underlying FaaS Engine creates the actual function runtime to be invoked by the Object Module.
@@ IV-D Object Module and FaaS Engine
Figure 5: The cluster of Invokers replicates and distributes object data across the cluster via IMDG with consistent hashing. The invoker offloads the invocation to a corresponding FaaS function.
@@
Recall that OaaS needs to support three types of functions: built-in, custom, and dataflow. Unlike built-in functions and dataflow functions that can be executed without the direct need of the FaaS engine, custom functions need to execute the developer-provided code on the FaaS engine. Thus, Oparaca requires a mechanism to utilize the FaaS engine to execute the custom function code while allowing it to access the object state transparently and with the minimum data transfer overhead. Needless to say, this mechanism also maintains the separation between the Object Module and the FaaS engine.
@@
To fulfill the above expectations, we design the object invocation mechanism in the Object Module by distinguishing between structured and unstructured states and managing it so that the data access overhead is minimized. We develop a hybrid approach that leverages the “pure function” technique for structured data access and the “datastore abstraction” technique for unstructured data access. The rationale of this design choice is that the unstructured state (i.e., BLOB) is usually large and expensive to transfer; hence, to maintain efficiency, the FaaS engine should retrieve the state directly from the object storage (e.g., S3) in a lazy, on-demand manner. This differs from the structured state, for which we include the state as an input argument to maintain a clear separation between the Object Module and the FaaS engine and let the FaaS engine maintain its statelessness.
@@
In the Oparaca architecture (Figure 4), he mechanism for handling invocation and state management is managed by the Invoker component. In particular, to offload the object invocation to the FaaS engine, Invoker bundles the request and the related structured object data as a “task”, as described in the next part, and passes it to the associated FaaS engine for execution.
IV-D1 Task Generation in the Invoker
@@
Upon receiving a function call, the Invoker bundles the invocation request and associated object data into the task and offloads it to be executed on the FaaS engine. To further reduce the data transfer overhead of providing the object abstraction in the task generation process, we design Invokers to maintain the object data (i.e., state and metadata) in a distributed hash table [20], thereby reducing the cost of data transfer in a scalable manner. As shown in Figure 5, we equip each Invoker instance with an embedded in-memory data grid (IMDG) [40]. IMDG partitions the entire data space into multiple segments and distributes them across Invoker instances. The Invoker with IMDG determines the segment for a given object by consistent hashing of the object ID and assigns the object data to the selected segment. Similarly, to retrieve the object data, IMDG determines the owner of the data and then fetches it from the owner of the segment in one hop.
@@ IV-D3 Task Completion
After the FaaS engine completes the task, it sends the task completion data to the Invoker to update the state. If the function reports a failed task, the state remains unchanged. Otherwise, the Invoker updates the object data in IMDG and then writes it to the persistent database immediately or asynchronously. If an invocation involves both structured and unstructured states, we use pure and datastore techniques together, which can potentially lead to “state inconsistency challenges”. We address this challenge in section IV-F.
@@ IV-D4 Synchronous and Asynchronous Invocation
As mentioned in Section IV-A and shown in Figure 5, we designed Oparaca to offer synchronous and asynchronous function invocations. In synchronous mode, the function is executed immediately upon invocation and returns the result to the caller. Meanwhile, in asynchronous mode, the invocation ID is provided to the caller as a reference so they can check the invocation result later. The request is placed into the message broker to be reliably processed at a later time. To accommodate both modes, the Invoker utilizes the handler instance to accept the invocation request for either the REST API (synchronous) or the message broker (asynchronous). Subsequently, the handler instance forwards the request to be processed in the same way by the other part of the Invoker.
@@ IV-F2 Resilience against race condition
Race conditions in Oparaca can occur when multiple invocations modify the same object data simultaneously, resulting in potential data inconsistency. One way to prevent this issue is by using database transactions; however, this method lacks abstraction as it allows direct function code access to the database and is tightly dependent on the type of database. An alternative approach to avoiding race conditions is the cluster-wide pessimistic locking mechanism to synchronize the locking state for all invokers. Nevertheless, this approach necessitates additional network communication to coordinate the locking state, which can lead to scalability issues. Alternatively, we develop an improved version of this mechanism, called “localized locking,” which relies on consistent hashing to direct the invocation request to the invoker that owns the primary copy of the targeted object data. Each invoker will only need to lock the object locally without additional network communication, making it more scalable than the cluster-wide version. Additionally, our localized locking approach guarantees that requests to the same object are executed in the arrival order, which is necessary in certain use cases where order matters, such as seat reservations. This is difficult to achieve with cluster-wide locking.
@@ IV-F3 Failure recovery in Oparaca
However, the message broker cannot completely address the third problem. That is, the Invoker can process the same invocation request more than once when the message broker has not acknowledged the completed one before the system failure occurs. We prevent this problem by tracking the offset number of the last processed request and adding it to each object metadata. In this manner, before processing an invocation request, Invoker checks the offset number of the target object to see if it is lower than the offset number of an incoming request. When the condition is met, the Invoker can detect that it has not been processed and perform the normal operation. Otherwise, it must be skipped to avoid reprocessing.
@@ IV-G Dataflow Abstraction in Oparaca
To offer a high-level abstraction to declare a workflow, Oparaca provides the dataflow abstraction as a built-in feature that enables developers to declaratively define the invocation steps as a directed acyclic graph (DAG) in a domain-specific language (DSL) with YAML format. In every step, the developer can declare the output of each invocation as a temporary variable within the workflow. Then, the next invocation can use the temporary variables from previous steps as the input or target to call the function. Upon registering a dataflow function by the developer, Invoker constructs the DAG by having the invocation step as the edge and the objects as nodes.
Upon calling the dataflow function, one of the Invokers takes on the role of orchestrator, similar to the orchestrator pattern [32] in microservices. It breaks down the dataflow into multiple lower-level invocations and forwards them based on the topological order of DAG. Using consistent hashing, the invoker can determine the address of the target object and send the request directly to another Invoker that holds the target object. When each step is completed, the orchestrator keeps track of the intermediate dataflow state to transparently operate the data exchange between invocation steps. With the orchestrator pattern, the dataflow control logic is centralized into a single invoker, simplifying the management, monitoring, and error-handling implementation.
When using the orchestrator pattern, the exact-once guarantee may be compromised because the object data is stored separately from the dataflow state. If the guarantee is needed, Oparaca allows flagging all invocation steps as immutable. Upon handling the dataflow request, Oparaca can generate the output ID in advance for each step, making each step of dataflow execution idempotent and safely re-executable.
@@ V-1 Security
Certain security measures can be implemented in Oparaca to strengthen it against potential attacks. The first measure is to reduce the attacking surface by limiting the necessary inbound traffic to the function container. As the function container is only accessed by the Invoker, the traffic policy can be configured to block inbound traffic except from the Invoker. The second measure is to avoid reusing secret tokens. To prevent the function container from accessing out-of-context data via analyzing the URL path, we use the presigned URL mechanism for object storage. Thus, object storage in Oparaca is more secure than in FaaS, where the same secret key is used for every request. To secure the storage adapter, we can make the Invoker generate a unique secret token for each task, and every request for the storage adapter must be authenticated via the secret token.
V-2 Multi-tenancy
The primary concern of multi-tenancy is ensuring data and resource isolation. The fundamental idea is to prevent sharing classes and functions among tenants. Since custom functions are offloaded and executed in a FaaS engine that provides strong isolation—with no shared functions—the execution environment is effectively contained within the FaaS engine. Regarding the Invoker and data management module, it is possible to share these components, as the data is stored separately in each class. However, depending on the billing model and isolation requirements, we can enhance security and resource isolation by separating these components for each tenant.
V-3 Cold Start
The developer functions and the Oparaca components can benefit from scale-to-zero to reduce the cost when there is no usage. However, this has the side effect of causing more cold starts. Since Oparaca components are shared across functions, we can effectively keep it warm to eliminate the additional cold start impact. In such a case, the cold start performance entirely depends on the underlying serverless execution engine.
@@ VI Performance Evaluation
Figure 7: The average execution time of functions for objects with various state sizes in synchronous and asynchronous invocations. Two versions of Oparaca are examined: the full version and the version without URL-redirection (oprc-relay). We also capture the time used by the internal Knative in both Oparaca versions and show them with the suffix -exec and plot them in the same bar as their Oparaca version.
@@ VI-A Experimental Setup
We deploy the Oparaca platform on 4 machines of Chameleon Cloud [25], each with 2 sockets of 24-Core Intel(R) Xeon(R) Gold 6240R CPU processors that collectively have 192 cores, 768 GB memory, and SSD SATA storage. We set up the Kubernetes cluster, which includes 15 VMs with 16 vCPUs and 32 GB of memory. We made another 2 VMs for the S3-compatible storage (Minio [21]) for unstructured data and ArangoDB ([2]) for structured data. Oparaca is implemented using Java with Infinispan [22] for IMDG.
@@
We used Gatling[13] for load generation and implemented three applications to serve as the workload. First is the video transcoding function, which utilizes FFmpeg [39], a CPU-intensive application. The second is a text concatenation function that concatenates the content of a text file (state) with an input string. This function represents a highly IO-intensive workload. Third is the JSON update function, which uses only structured data in JSON and is used to insert key-value pairs into the JSON state data randomly. The remaining workload characteristics are specific to each experiment and are explained in the respective sections. All three functions are implemented in the Python language.
@@ VI-B Analyzing the Imposed Overhead of Oparaca
For the transcoding function (Figures 8a and 8d), at the low concurrency levels (<< 80 invocations), Oparaca has average response times higher than Knative, but for the higher concurrency levels, the response time of Knative grows faster than Oparaca due to computing resource limitations. Oparaca doesn’t need to fetch video file metadata, giving it an edge at high concurrency. In the concatenation function (Figures 8b and 8e), however, this phenomenon does not happen. The difference is that text concatenation is IO-intensive and desires high network bandwidth. The overhead of unstructured data access overwhelms the performance gain from eliminating structured data fetching.
@@ VI-C Scalability of the Oparaca Platform
According to Figure 9a, the speedup of Knative plateaus after reaching 6 VMs. We realized that this plateau is attributed to the database write operation throughput bottleneck. Conversely, Oparaca exhibits the potential for higher speedup enhancement due to its reliance on the distributed in-memory hash table to consolidate data for batch write operations. This approach can boost maximum throughput by up to 3.27x when comparing oprc-bypass with knative.
@@ VI-D Performance of Localized Locking
To analyze effectiveness of localized locking, we created a variation of Oparaca, called oprc-cl, that has cluster-wide locking and evaluated it using a cluster of 12 Invokers while increasing the request arrival rate to measure the locking overhead. To generate requests involving the locking mechanism, we created multiple requests targeting the same object. From Figure 10, the overhead of localized locking remains mostly constant, while the overhead of cluster-wide locking rises for higher request rates. The cluster-wide version does not exhibit this behavior, as the network communication overhead limits the throughput and hinders high invocation rates.